You are sitting in a plane, in straight horizontal flight at constant speed of 250 \text{m.s}^{-1}m.s 

−1

 , when you drop your pen from height 0.30 m above your leg. You've just shown that it hits your leg after 0.25 s.


1
Expert's answer
2020-09-29T09:43:00-0400
"h=0.5gt^2\\t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(0.30)}{9.8}}=0.25\ s"

The answer is right!